OpenStudy (anonymous):

Write the solution in interval notation: x < 6 and x >= -3

OpenStudy (anonymous):

-3<x<6 so x can by anywhere from between -3 to 6

OpenStudy (anonymous):

\[-3 \le x < 6\]

OpenStudy (saifoo.khan):

it can be written in many other ways too, which do u want???

OpenStudy (anonymous):

\[(-\infty, 6) intersection [-3, +\infty)\]

OpenStudy (saifoo.khan):

hmm,.

OpenStudy (anonymous):

[-3, 6)

OpenStudy (anonymous):

Thanks, I figured this one out on my own. :)

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

=)